OpenStudy (anonymous):

Just need verification! We do not know the equation of y=f(x), but we do know that f'(x)=4/3 x^(2/3) (x-1). Need the critical values. I calculated them to be 0 and 1. Can anyone verify that I am on track?

OpenStudy (anonymous):

Syntax is a touch difficult to understand but I think you mean \[f'(x)= \frac{ 4 }{ 3 }(x-1)x^\frac{ 2 }{ 3 }\] In which case the critical values obviously being the values that turn the derivative to 0, in which case yes, they would be 0 and 1 :)

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } x ^ (2/3) (x-1)\]

OpenStudy (anonymous):

The power to the x is 2/3.

OpenStudy (anonymous):

Would this still give critical values of 0 and 1?

OpenStudy (anonymous):

Yes it would :)